OpenStudy (anonymous):

Complete the square for the binomial. Then factor the resulting perfect square trinomial. x2 + 10x A) x2 + 10x + 25 = (x + 5)2 B) x2 + 10x + 5 = (x + 25)2 C) x2 + 10x + 10 = (x + 100)2 D) x2 + 10x + 100 = (x + 10)2

hero (hero):

A

OpenStudy (anonymous):

\[(1x)^2+2\cdot1\cdot5+(5)^2=x^2+10x+25=(x+5)^2\]